# bibliography.bib

### v2

@@ 5.2 Elliptic PDE with lognormal coefficients – direct solver
In this section, we again solve (27) with the same input data, i.e., $f\equiv 1$, $s=4$, $q=2$, and $\sigma=2$. The coarsest mesh size is again $h_{0}=1/8$. However, here the underlying linear system (with symmetric positive definite matrix) is solved using a double precision Cholesky factorisation from PETSc in the standard MLMC algorithm and a low precision Cholesky factorisation with iterative refinement in our MPML algorithm. Recall that throughout our experiments by standard MLMC we mean Algorithm 4.1 without step 2, where the computations in step 4 are (here) carried out with double precision. We use our own implementation of iterative refinement and of the low precision Cholesky factorisation. To carry out computations in half, single, and double precision, we use the numerical types of NumPy, namely float16(), float32(), and float64() for half, single, and double precision, respectively. Apart from the linear solver, all calculations are carried out in double precision.
@@
We continue by estimating the cost gains using MPML estimator $\widehat{Q}_{L,\{N_{l}\}}^{\text{MPML}}$ with low precision Cholesky factorisation and iterative refinement compared to standard MLMC estimator $\widehat{Q}_{L,\{N_{l}\}}^{\text{ML}}$ with double precision Cholesky. Again, both estimators use the same numbers of samples $\{N_{l}\}$. Since memory references are by far the most costly part on modern computing architectures, both in terms of time and energy cost (cf. the discussion Section 2.4), we use memory access as a simple cost model, in terms of total number of bits loaded into main memory. Thus, accessing a half precision floating point number is 2×2$\times$2 × cheaper than accessing a single precision number and 4×4$\times$4 × cheaper than accessing a double precision number. The simulated cost gain can then be estimated by counting the number of entries in all vectors and the number of non-zeros in all sparse factorisations of the matrices computed and stored in the process of solving the linear systems during the iterative refinement process. Figure 6 shows the total cost gain per level (cost per sample $\times$ number of samples) for $\widehat{Q}_{L,\{N_{l}\}}^{\text{MPML}}$ compared to $\widehat{Q}_{L,\{N_{l}\}}^{\text{ML}}$. On all levels we observe a simulated memory gain of $\approx 2$. The total memory gain (sum over all levels) is $\approx 2$ as well.
Figure 7: Comparison of estimated MSE for MPML and MLMC using the same number of samples and the same random seeds (with approximate $95\%$ confidence intervals). MLMC uses double precision Cholesky, while MPML uses low precision Cholesky with iterative refinement (here with quarter precision on level 00).
Figure 8: Comparison of total costs of the estimators per level in terms of memory access for MPML and MLMC. MLMC uses double precision Cholesky, while MPML uses low precision Cholesky with iterative refinement (here with quarter precision on level 00).
In the last example, we show how further memory gains can be obtained using quarter precision on the coarsest level. For this experiment, we again solve (27) with $f\equiv 1$ on $D$, $s=1$ and $\sigma=1$, but now we choose the coarsest mesh size to be $h_{0}=1/4$. To determine the effective precision $\delta_{l}$ in (24) we choose $k_{p}=0.4$. Iterative refinement with the Cholesky solver is used with the following settings: $qhhh$ on level 00, and $hhss$ on levels $1$ to $3$. This means that on level 00 where $h_{0}=1/4$, quarter precision (q43, see Section 2.1) is used for the Cholesky factorisation. To simulate quarter precision, the pychop package of Xinye Chen was used; see [pychop]. Figures 7 and 8 show the MSE for MPML and MLMC and the simulated memory gain, respectively. They have been produced analogously to Figures 5 and 6. We observe that for the MSE tolerance $2\times 10^{-6}$, the MPML relative error is $11\%$ bigger compared to standard MLMC error with the same number of samples, while we are able to achieve a total memory gain of $\approx 3.5$. Recall that the cost gain is simulated by counting non-zeros computed and stored in the process of solving the linear system $Ax=b$.
We have not encountered overflow in any of the examples presented in this section when working with the lower precisions. It is important to note that due to the fact that the coefficients of the PDE are lognormally distributed, overflow can occur with low probability. In the case where overflow occurs, scaling or shifting techniques can be used; see [higham2019squeezing].
@@ 6 Conclusion
Through two practical examples, we showcased the potential of our approach to obtain significant cost gains. Using a low-precision sparse direct solver with iterative refinement, we achieved a simulated memory gain of up to 3.5×3.5$\times$3.5 ×, while employing a MINRES iterative solver yielded a speedup of 2×2$\times$2 × in floating point operations. These results highlight the significant potential for energy-aware scientific computing. The potential for future work lies in generalising this approach to other uncertainty quantification frameworks, such as multilevel Markov chain Monte Carlo or the multilevel stochastic collocation method and exploring its broader applications in scientific computing. The efficient parallel implementation of low-precision solvers within multilevel sampling methods is of great interest and promises to bring further speedup.

### v3

@@ 5.2 Elliptic PDE with lognormal coefficients – direct solver
In this section, we again solve (29) with the same input data, i.e., $f\equiv 1$ on $D=(0,1)^{2}$, $s=4$, $q=2$, and $\sigma=2$. The coarsest mesh size is again $h_{0}=1/8$. However, here the underlying linear system (with symmetric positive definite matrix) is solved using a double precision Cholesky factorisation from PETSc in the standard MLMC algorithm and a low precision Cholesky factorisation with iterative refinement in our MPML algorithm. Recall that throughout our experiments by standard MLMC we mean Algorithm 4.1 without step 2, where the computations in step 4 are (here) carried out with double precision. We use our own implementation of iterative refinement and of the low precision Cholesky factorisation. To carry out computations in half, single, and double precision, we use the numerical types of NumPy, namely float16(), float32(), and float64() for half, single, and double precision, respectively. Apart from the linear solver, all calculations are carried out in double precision.
@@
We continue by estimating the cost gains using MPML estimator $\widehat{Q}_{L,\{N_{l}\}}^{\text{MPML}}$ with low precision Cholesky factorisation and iterative refinement compared to standard MLMC estimator $\widehat{Q}_{L,\{N_{l}\}}^{\text{ML}}$ with double precision Cholesky. Again, both estimators use the same numbers of samples $\{N_{l}\}$. Since memory references are by far the most costly part on modern computing architectures, both in terms of time and energy cost (cf. the discussion Section 5.3), we use memory access as a simple cost model, in terms of total number of bits loaded into main memory. Thus, accessing a half precision floating point number is $2\times$ cheaper than accessing a single precision number and $4\times$ cheaper than accessing a double precision number. The simulated cost gain can then be estimated by counting the number of entries in all vectors and the number of non-zeros in all sparse factorisations of the matrices computed and stored in the process of solving the linear systems during the iterative refinement process. Figure 7 shows the total cost gain per level (cost per sample $\times$ number of samples) for $\widehat{Q}_{L,\{N_{l}\}}^{\text{MPML}}$ compared to $\widehat{Q}_{L,\{N_{l}\}}^{\text{ML}}$. On all levels we observe a simulated memory gain of $\approx 2$. The total memory gain (sum over all levels) is $\approx 2$ as well.
In the last example, we show how further memory gains can be obtained using quarter precision on the coarsest level. For this experiment, we again solve (29) with $f\equiv 1$ on $D$, $s=1$ and $\sigma=1$, but now we choose the coarsest mesh size to be $h_{0}=1/4$. To determine the effective precision $\delta_{l}$ in (16) we choose $k_{p}=0.4$. Iterative refinement with the Cholesky solver is used with the following settings: qhhhqhhh on level 0, and hhsshhss on levels $1$ to $3$. This means that on level 0 where $h_{0}=1/4$, quarter precision (q43, see Section 2.1) is used for the Cholesky factorisation. To simulate quarter precision, the pychop package of Xinye Chen was used; see [7]. Figures 8 and 9 show the MSE for MPML and MLMC and the simulated memory gain, respectively. They have been produced analogously to Figures 6 and 7. We observe that for the MSE tolerance $2\times 10^{-6}$, the MPML relative error is $11\%$ bigger compared to standard MLMC error with the same number of samples, while we are able to achieve a total memory gain of $\approx 3.5$. Recall that the cost gain is simulated by counting non-zeros computed and stored in the process of solving the linear system $Ax=b$.
We have not encountered overflow in any of the examples presented in this section when working with the lower precisions. It is important to note that due to the fact that the coefficients of the PDE are lognormally distributed, overflow can occur with low probability. In the case where overflow occurs, scaling or shifting techniques can be used; see [22].
@@
Figure 8: Comparison of estimated MSE for MPML and MLMC using the same number of samples and the same random seeds (with approximate $95\%$ confidence intervals). MLMC uses double precision Cholesky, while MPML uses low precision Cholesky with iterative refinement (here with quarter precision on level 0).
Figure 9: Comparison of total costs of the estimators per level in terms of memory access for MPML and MLMC. MLMC uses double precision Cholesky, while MPML uses low precision Cholesky with iterative refinement (here with quarter precision on level 0).
5.3 Energy savings through iterative refinement
We obtained a simulated gain of up to $3.5\times$ in terms of memory references in our experiments, see Figure 7. The ratio of the cost of one memory access to one floating point operation is continuing to grow in recent architectures [25, Table 2]. On modern 7nm semiconductor architectures, a cache (SRAM) memory access is on average $10\times$ to $100\times$ more expensive in terms of energy than a floating point operation, while accessing SRAM is between $10\times$ and $100\times$ cheaper than accessing DRAM. Therefore, an efficient parallel implementation of iterative refinement within the proposed adaptive algorithm (Algorithm 4.1) that allows one to store more data in the cache memory can bring a significant reduction in energy cost, which is a promising area for future research.
@@ 6 Conclusion
Through two practical examples, we showcased the potential of our approach to obtain significant cost gains. Using a low-precision sparse direct solver with iterative refinement, we achieved a simulated memory gain of up to $3.5\times$, while employing a MINRES iterative solver yielded a speedup of $2\times$ in floating point operations. While these results highlight the significant potential for energy-aware scientific computing, the presented examples serve as a proof-of-concept for the developed method rather than a robust demonstration across a spectrum of problem difficulties. The potential for future work lies in applying this approach to other uncertainty quantification frameworks, such as multilevel Markov chain Monte Carlo or the multilevel stochastic collocation method and exploring its broader applications in scientific computing. The efficient parallel implementation of low-precision solvers within multilevel sampling methods is of great interest and promises to bring further speedup.